A particle executes simple harmonic motion
with an amplitude of 2.69 cm.
At what positive displacement from the
midpoint of its motion does its speed equal
one half of its maximum speed?
Answer in units of cm.
1
Expert's answer
2020-02-17T09:16:28-0500

From the conservation of energy:


"0.5mV^2=0.5kA^2=0.5m(0.5V)^2+0.5kx^2"

"0.5kx^2=0.5mV^2-0.5m(0.5V)^2=(0.75)0.5mV^2=(0.75)0.5kA^2"

"x=\frac{\sqrt{3}}{2}A=\frac{\sqrt{3}}{2}2.69=2.33\ cm"